# Coward: Collision-based Watermark for Proactive Federated Backdoor Detection

### v3

@@ 3 Revisiting Existing Detection Methods and Multi-Backdoor Interplay
In this section, we first introduce the threat model and revisit the limitations of existing detection methods. We then examine the interplay effects of multi-backdoor attack, which directly motivate our detection method.
@@ 3.1 Preliminary: Threat Model
We define the threat model by specifying the goals and capabilities of both the attacker and the defender.
Attacker’s Goal and Capability. The attacker resides on the client side and aims to manipulate the federated system by injecting a targeted backdoor into the global model, causing it to misclassify all trigger-embedded inputs into a designated target class while maintaining normal performance on benign inputs. Although the attacker cannot directly access the global model, they may control a subset of clients and submit arbitrary malicious models or gradients to the server, thereby influencing the global model updates.
@@ 3.2 Revisiting Passive Detection
Many existing passive detection methods assumed that malicious updates deviated from benign updates in the parameter space due to the additional optimization of the backdoor objective during local training [blanchard2017machine]. Based on this assumption, many methods leveraging clustering and outlier filtering were proposed to detect malicious clients [RFA, nguyen2022flame]. To validate this assumption under practical non-i.i.d. settings, we analyzed the statistical characteristics of local model updates, and show its correlation to misjudgements of benign clients.
Settings. We revisit gradient norm distribution for each client, under a highly non-i.i.d. partitioned CIFAR-10 dataset with label distribution generated by Dirichlet prior with $\alpha=0.3$, and 10 out of 100 random clients participation. One malicious client performs backdoor attack using a pixel-style trigger targeting class 0.
@@
(b) Norm-Misjudgement Correlation Analysis
@@
Results. As shown in Figure 2(a), the suspicion toward attack is significantly reduced due to the severity of data heterogeneity: (1) in the first column, we present the size and label skewness of local datasets. The blue line indicates the dataset size, while the bar height represents the skewness, measured by the standard deviation of class-wise sample counts. We can observe that client data distributions are highly diverse. Notably, malicious clients do not exhibit abnormal characteristics in either size or skewness; their distributions fall within a typical range. In contrast, some benign clients show extreme distributions with both large data volumes and high skewness. (2) In the second column, we plot the mean and standard deviation of gradient norms uploaded by all 100 clients across training rounds, sorted by their mean values. The results reveal that the malicious client’ updates have slightly smaller gradient norms compared to most benign clients and appear inconspicuous amid the naturally diverse benign client updates. (3) In the third column, we visualize the top-2 PCA component of local and global models (based on the last-layer weights) during training. The results show that the backdoor models’ update direction and distance lie in a moderate region, appearing unharmful due to the diverse update directions introduced by the non-i.i.d. distribution. Figure 2(b) further presents a correlation analysis, revealing a clear association between model update norms and misjudgement of malicious clients. Specifically, the x-axis shows the mean gradient norm, the y-axis indicates the number of misjudgments ($i.e.$, times a benign client was incorrectly flagged), and the bubble size reflects the standard deviation of the gradient norm. The evident upward trend suggests that clients with larger and more variable gradient norms are more susceptible to be misjudged. In conclusion, we have:
Takeaway 1: Data heterogeneity ($i.e.$, non-i.i.d.) amplify client model variance, blurring distinctions between malicious and benign clients, diminishing the effectiveness of passive backdoor detection.
To mitigate this limitation, proactive methods decouple the detection reliance from parameter differences to client-side reactions against server-side intervention actions, which effectively alleviate sensitivity to data heterogeneity.
@@
Figure 3: The misdirection effect of OOD bias against the existing proactive detection methods. With the attacker-specified target class set to ‘0’ (red), we show five clients from the same round, presenting their OOD prediciton distribution, with results in no-planting case the first row and planting case in the second row. The gray subfigure shows the local data distribution. The red dashed line marks the detection threshold; classes with inspection accuracy above this line are flagged as malicious and highlighted in red border. The results indicate that: (1) Even without server-side planting, non-target classes may exhibit OOD-induced high inspection accuracy (orange) on benign clients; (2) With planting, misjudgment increases, and more classes tend to exhibit higher inspection accuracy.
@@
Figure 4: Exacerbation of OOD bias induced by existing proactive detection methods. The Indicator method intensifies both the magnitude and variation of OOD bias.
@@ 3.3 Revisiting Proactive Detection
As a pioneering work in proactive detection, Indicator identified a co-existing ($i.e.$, maintaining) effect, where a subsequently planted backdoor with the same target label could preserve the effect of a previously injected one. However, due to the privacy-preserving constraint of federated learning, it could not access client data and had to rely solely on OOD data to modify the global model. To address this constraint, Indicator implanted random OOD mappings ($i.e.$, assigning labels to OOD samples via uniform random sampling) into the global model to activate the co-existing effect. During detection, it evaluated each client’s class-wise prediction accuracy on these OOD data: any class exceeding a predefined threshold flagged the client as malicious, and the corresponding class was assumed to be the attacker’s backdoor target. Such a detection mechanism no longer relies on parameter space, making it less susceptible to the distraction effect of non-i.i.d. data and thus more effective than passive detection methods.
@@
However, leveraging OOD samples to activate the co-existing effect introduces a vulnerability to OOD prediction bias, where deep neural networks inevitably tend to classify OOD samples into arbitrary classes with high confidence [nguyen2015ood_bias, guo2017ood_bias_calib, hein2019relu_ood_bias_theory]. Specifically, it leads to a high false positive rate, as analyzed in the following parts.
@@
Settings. We analyze the prediction distribution of OOD samples under Indicator [li2024backdoorindicator], both with and without server-side intervention. The main FL task is CIFAR-10 classification, with EMNIST used for proactive pattern planting. During planting, Indicator randomly partitioned OOD samples in the label space with same 10% ratio. In each round, 10 clients are sampled from a total of 100, with one attacker targeting class 0. The data heterogeneity is controlled by a Dirichlet distribution with $\alpha=0.3$.
@@
Results. As shown in Figure 3, we find that: (1) under the no-intervention setting, for the malicious client, OOD samples successfully activate the learned backdoor pattern with high accuracy, revealing an empirical correlation between the backdoor trigger and the OOD pattern. Meanwhile, for benign clients, certain classes exhibit inspection accuracy that approaches or even exceeds the predefined detection threshold of 95%, resulting in significant misjudgment; (2) Under the with-intervention setting, the pre-planted OOD pattern leads to similar effects and, in some cases, causes even more classes keep higher accuracy among benign clients ($i.e.$, tending to be treated as malicious clients). These phenomena indicate that:
@@
Takeaway 2: OOD prediction bias is an inherent vulnerability of DNNs [nguyen2015ood_bias, guo2017ood_bias_calib, hein2019relu_ood_bias_theory], where OOD samples tend to be classified with high confidence into arbitrary classes. This causes benign clients to exhibit unexpectedly high inspection accuracy during Indicator’s detection phase, leading to frequent yet unpredictable false positives.
@@
Going further, we hereby demonstrate how the intervention action intensifies OOD prediction bias in Figure 4, from which we draw the following key observations. (1) Disruption of temporal consistency: In subfigure (a), we present the prediction distributions of a specific client across multiple global rounds. The no-intervention setting maintains a stable bias toward a particular class over time, whereas the with-intervention setting exhibits a highly variable trend, $i.e.$, the class with the highest prediction accuracy fluctuates significantly across rounds. This instability is likely caused by the random label assignment used in the Indicator method, where OOD samples with similar semantic content may be mapped to different labels during the intervened training process, creating an disordered association between the OOD space and the task label space; (2) Amplified bias level and variation: As shown in subfigure (b), we quantify the overall bias level and its temporal variation across all 100 clients, sorted by bias magnitude. The results show that Indicator’s intervention introduces significantly higher bias intensity and instability compared to the no-intervention case. We further analyze the correlation between OOD bias levels and false positive rates in Section 5(Figure 7), showing a strong positive relationship. Building on above evidence, we arrive at the conclusion that:
@@
Given the limitations of random OOD planting, the effectiveness of leveraging the behind multi-backdoor coexistence effect introduced by Indicator becomes questionable. This naturally leads to a deeper question: Does an alternative multi-backdoor interaction mechanism exist, and if so, can it offer a more robust foundation for utilizing OOD data? Accordingly, we shift to revisit multi-backdoor interplay.
@@ 3.4 Revisiting Multi-Backdoor Interplay
TABLE I: Types of multi-backdoor interplay.
Different Target Same Target
Different Trigger collision (w./ BN) coexistence (w./ BN)
Same Trigger collision coexistence
@@
The Demonstration of Collision Effect. We conduct the experiment under a centralized setting simulating two entity sequentially plant backdoors into a same ResNet model on CIFAR-10 dataset but with different target label. Specifically, the first backdoor uses a WaNet trigger targeting class 1 while the second backdoor uses a pixel-pattern trigger targeting class 0. Following [li2024backdoorindicator], the batch normalization statistics are also switched accordingly. As shown in Figure 5, the left subfigure demonstrates that the ASR of the first backdoor steadily decreases as the second backdoor is progressively injected. In contrast, the right subfigure shows that the ASR decline from normal fine-tuning (due to knowledge forgetting) remains clearly less severe, highlighting the distinct collision effect between the two backdoors.
@@
Collision Effect v.s. Coexistence Effect. To this end, we can get a complete understanding of multi-backdoor interplay under sequential implantation: Backdoors targeting the same class tend to coexist, while those with different targets often interfere with each other. To be comprehensive, we summarize all interaction cases in Table I, where “w./ BN” denotes for “with BN switch”. Though seemingly opposite, the two mechanisms can be understood through a unified and intuitive perspective. As noted by Indicator [li2024backdoorindicator], backdoor triggers can be interpreted as OOD features in the benign latent space, typically residing in a shared OOD region within the latent space. When multiple triggers share the same target label, they construct consistent mappings from the latent OOD region to that label. This alignment allows backdoors to reinforce each other, resulting in their coexistence. In contrast, when triggers are associated with different target labels, the newly implanted trigger disrupts the existing mapping from the latent OOD region to the prior label, leading to collision or substitution.
@@
Insights from Collision Effect. Based on the collision effect, it becomes feasible to design a backdoor-based watermark that intentionally interferes with the attacker’s trigger while remaining intact in benign clients. This leads to a new proactive detection paradigm distinct from the existing coexistence-based methods, where client models that fail to retain the watermark ($i.e.$, exhibit low inspection accuracy) are identified as malicious, while those that preserve it ($i.e.$, show high inspection accuracy) are considered benign. Under this inverted paradigm, the high inspection accuracy induced by OOD prediction bias in benign clients no longer misleads detection, instead, it becomes beneficial in some cases. Building on these insights, we propose our collision-based watermarking method, as detailed in the next section.
@@
Figure 5: The collision effect between target-different backdoors. (Left): A subsequently injected backdoor significantly degrades the ASR of the first backdoor. (Right): The ASR degradation of the first backdoor caused by standard fine-tuning is notably smaller than that caused by subsequent backdoor injection. Together, they clearly demonstrate the presence of a collision effect.
@@ 4 Methodology
Figure 6 shows an overview of our approach, consisting of three stages: watermark injection, watermark interaction, and watermark detection. We describe each in detail below.
@@ 4.1 Watermark Injection
Watermarking is common for integrity protection by embedding identifiable patterns into models or data [liu2024wm_llm_survey, wei2024wm_tifs24_yml, qu2024provably]. Inspired by this idea and our observation of the backdoor collision effect, we leverage a backdoor-based watermarking strategy to identify malicious clients in FL. To adapt the idea of backdoor watermarking to the out-of-distribution setting, we first revisit the conceptual structure of traditional backdoor mechanisms. In conventional in-distribution scenarios, a backdoor can be viewed as composed of two parts: (1) a standard task mapping that align clean inputs to its ground-truth label, and (2) a tampered mapping that redirects triggered inputs to attacker-specified labels. Following this intuition, we borrow a similar two-part structure in the OOD context: an OOD base mapping for clean OOD samples and an OOD watermark mapping for triggered ones.
@@
Planting a Regulated Base Mapping. In FL, the OOD planting set and client training data share the same global model, thus operating within the same prediction space. However, their ground-truth semantic spaces are inherently different. This misalignment introduces prediction ambiguity, where the same output value may imply conflicting meanings. For instance, label “0” might refer to the digit “0” in an EMNIST-based OOD planting set, but to “airplane” in a CIFAR-10 client task. Such ambiguity can lead to two negative effects: (1) the OOD mapping may distort client-learned semantics, reducing benign accuracy; (2) the non-i.i.d. nature of client training may corrupt the OOD mapping.
To address these problems, we manually assign a fix one-to-one label mapping $\pi_{1}$ from the OOD label space $\mathcal{Y}_{o}$ to the task label space $\mathcal{Y}_{t}$. This defines the base mapping planting:
$\pi_{1}:\mathcal{Y}_{o}\rightarrow\mathcal{Y}_{t},$ (2)
$\mathcal{D}_{p}=\{(\bm{x}_{i},\pi_{1}(y_{i}))\mid y_{i}\sim\mathcal{Y}_{o}\},$ (3)
$\mathcal{L}_{base}=\sum_{i\in\mathcal{D}_{p}}\mathrm{CE}(f(\bm{x}_{i};\bm{w}),\pi_{1}(y_{i})),$ (4)
@@
Here, $\mathcal{D}_{p}$ denotes the planting set, $\bm{w}$ represents the model parameters, and $\mathrm{CE}$ is the cross-entropy loss. Since OOD samples lack explicit semantic grounding in the task label space, the mapping’s meaning is not essential, what matters is that it remains fixed. Our choice of consistent one-to-one assignment stabilizes the ambiguity, reducing uncertainty in both OOD and task predictions. This contrasts with the design in BackdoorIndicator, which adopts random and dynamic base mappings, leading to greater instability. Instead, our method emphasizes a controlled and interpretable structure, as supported by the empirical results in Figure 7.
@@
Planting a Targeted Watermark Mapping. Building upon the established base OOD mapping, we further implant a targeted watermark mapping, analogous to the all-to-one backdoor setting [ba_survey_yiming, chameleon, yang2024_ba_atk] in centralized learning:
$\pi_{2}:\mathcal{Y}_{o}\rightarrow y_{m},y_{m}\in\mathcal{Y}_{t},$ (5)
$\mathcal{D}_{w}=\{(\text{tri}(\bm{x}_{i},\bm{t}),\pi_{2}(y_{i}))\mid(\bm{x}_{i},y_{i})\sim\mathcal{D}_{p}\},$ (6)
@@
Following prior works [gan2023towards, li2024backdoorindicator], we save and restore the running mean and variance of BatchNorm layers before and after watermark planting. This preserves the distinct feature distributions of the task and OOD spaces during training. In addition, to prevent the planted watermark from distorting the task-space sample-label mapping, we also apply a global model regularization term. Together, we get the overall loss:

### v4

@@ 3 Revisiting Existing Detection Methods and Multi-Backdoor Interplay
In this section, we first introduce the threat model and revisit the limitations of existing detection methods. We then examine the interplay effects of multi-backdoor attacks, which directly motivate our detection method.
@@ 3.1 Preliminary: Threat Model
Attacker’s Goal and Capability. The attacker resides on the client side and aims to inject a targeted backdoor into the global model, causing it to misclassify all trigger-embedded inputs into a designated target class while maintaining normal performance on benign inputs. Although the attacker cannot directly access the global model, they may control a subset of clients and submit arbitrary malicious models or gradients to the server, thereby influencing the global model updates.
@@ 3.2 Revisiting Passive Detection
Many existing passive detection methods assume that malicious updates deviate from benign updates in the parameter space due to the additional optimization of the backdoor objective during local training [4]. Based on this assumption, many methods leveraging clustering and outlier filtering have been proposed to detect malicious clients [31, 27]. To validate this assumption under practical non-i.i.d. settings, we analyze the statistical characteristics of local model updates and show their correlation with misjudgments of benign clients.
Settings. We revisit the gradient norm distribution for each client under a highly non-i.i.d. partition of the CIFAR10 dataset, with label distribution generated by a Dirichlet prior with $\alpha=0.3$, and participation by 10 randomly clients out of 100. One malicious client performs a backdoor attack using a pixel-style trigger targeting class 0.
@@
(b) Norm-Misjudgment Correlation Analysis
@@
Results. As shown in Figure 2(a), the suspicion toward attack is significantly reduced due to the severity of data heterogeneity: (1) in the first column, we present the size and label skewness of local datasets. The blue line indicates the dataset size, while the bar height represents the skewness, measured by the standard deviation of class-wise sample counts. We can observe that client data distributions are highly diverse. Notably, malicious clients do not exhibit abnormal characteristics in either size or skewness; their distributions fall within a typical range. In contrast, some benign clients show extreme distributions with both large data volumes and high skewness. (2) In the second column, we plot the mean and standard deviation of gradient norms uploaded by all 100 clients across training rounds, sorted by their mean values. The results reveal that the malicious client’s updates have slightly smaller gradient norms compared to most benign clients and appear inconspicuous amid the naturally diverse benign client updates. (3) In the third column, we visualize the top-2 PCA component of local and global models (based on the last-layer weights) during training. The results show that the backdoor models’ update direction and distance lie in a moderate region, appearing unharmful due to the diverse update directions introduced by the non-i.i.d. distribution. Figure 2(b) further presents a correlation analysis, revealing a clear association between model update norms and misjudgment of malicious clients. Specifically, the x-axis shows the mean gradient norm, the y-axis indicates the number of misjudgments, and the bubble size reflects the standard deviation of the gradient norm. The upward trend suggests that clients with larger and more variable gradient norms are more susceptible to being misjudged. In conclusion:
Takeaway 1: Data heterogeneity ($i.e.$, non-i.i.d.) amplifies client model variance, blurring distinctions between malicious and benign clients and diminishing the effectiveness of passive backdoor detection.
To mitigate this limitation, proactive methods decouple detection from parameter differences and instead rely on client-side reactions to server-side intervention actions, which effectively alleviates sensitivity to data heterogeneity.
@@
Figure 3: The misdirection effect of OOD bias against the existing proactive detection methods. With the attacker-specified target class set to ‘0’ (red), we show 5 clients from the same round, presenting their OOD prediction distribution, with results in the no-planting case in the first row and the planting case in the second row. The gray subfigure shows the local data distribution. The red dashed line marks the detection threshold; classes with inspection accuracy above this line are flagged as malicious and highlighted in red border. The results indicate that: (1) Even without server-side planting, non-target classes may exhibit OOD-induced high inspection accuracy (orange) on benign clients; (2) With planting, misjudgment increases, and more classes exhibit higher inspection accuracy.
@@
Figure 4: Exacerbation of OOD bias induced by existing proactive methods. The Indicator method intensifies both the magnitude and variation of OOD bias.
@@ 3.3 Revisiting Proactive Detection
As a pioneering work in proactive detection, Indicator identified a co-existing ($i.e.$, maintaining) effect, where a subsequently planted backdoor with the same target label could preserve the effect of a previously injected one. However, due to the privacy constraint of federated learning, it could not access client data and had to rely solely on OOD data to modify the global model. To address this constraint, Indicator implanted random OOD mappings ($i.e.$, assigning labels to OOD samples via uniform random sampling) into the global model to activate the co-existing effect. During detection, it evaluated each client’s class-wise prediction accuracy on these OOD data: any class exceeding a predefined threshold flagged the client as malicious, and the corresponding class was assumed to be the attacker’s backdoor target. Such a detection mechanism no longer relies on parameter space, making it less susceptible to the distraction effect of non-i.i.d. data and thus more effective than passive detection methods.
@@
However, leveraging OOD samples to activate the co-existing effect introduces a vulnerability to OOD prediction bias, where deep neural networks inevitably tend to classify OOD samples into arbitrary classes with high confidence [26, 15]. Specifically, it leads to a high false positive rate, as analyzed in the following parts.
@@
Settings. We analyze the prediction distribution of OOD samples under Indicator [17], both with and without server-side intervention. The main FL task is CIFAR-10 classification, with EMNIST used for proactive pattern planting. During planting, Indicator randomly partitioned OOD samples in the label space with the same 10% ratio. In each round, 10 clients are sampled from a total of 100, with one attacker targeting class 0. The data heterogeneity is controlled by a Dirichlet distribution with $\alpha=0.3$.
@@
Results. As shown in Figure 3, we find that: (1) under the no-intervention setting, for the malicious client, OOD samples successfully activate the learned backdoor pattern with high accuracy, revealing an empirical correlation between the backdoor trigger and the OOD pattern. Meanwhile, for benign clients, certain classes exhibit inspection accuracy that approaches or even exceeds the predefined detection threshold of 95%, resulting in significant misjudgment; (2) Under the with-intervention setting, the pre-planted OOD pattern leads to similar effects and, in some cases, causes even more classes to maintain higher accuracy among benign clients ($i.e.$, tending to be treated as malicious clients). These phenomena indicate that:
@@
Takeaway 2: OOD prediction bias is an inherent vulnerability of DNNs [26, 15], causing OOD samples to be classified into arbitrary classes with high confidence. Under Indicator’s detection scheme, this bias can give benign clients unexpectedly high suspicion scores, leading to frequent yet unpredictable false positives.
@@
Going further, we hereby demonstrate how the intervention action intensifies OOD prediction bias in Figure 4, from which we draw the following key observations. (1) Disruption of temporal consistency: In subfigure (a), we present the prediction distributions of a specific client across multiple global rounds. The no-intervention setting maintains a stable bias toward a particular class over time, whereas the with-intervention setting exhibits a highly variable trend, $i.e.$, the class with the highest prediction accuracy fluctuates significantly across rounds. This instability is likely caused by the random label assignment used in the Indicator method, where OOD samples with similar semantic content may be mapped to different labels during the intervened training process, creating a disordered association between the OOD space and the task label space; (2) Amplified bias level and variation: As shown in subfigure (b), we quantify the overall bias level and its temporal variation across all 100 clients, sorted by bias magnitude. The results show that Indicator’s intervention introduces significantly higher bias intensity and instability compared to the no-intervention case. We further analyze the correlation between OOD bias levels and false positive rates in Section 5(Figure 7), showing a strong correlation. To this end, we conclude that:
@@
Given the limitations of random OOD planting, the effectiveness of leveraging the multi-backdoor coexistence effect behind Indicator becomes questionable. This naturally leads to a deeper question: Does an alternative multi-backdoor interaction mechanism exist, and if so, can it offer a more robust foundation for utilizing OOD data? Accordingly, we shift to revisit multi-backdoor interplay.
@@ 3.4 Revisiting Multi-Backdoor Interplay
The Demonstration of Collision Effect. We conduct the experiment under a centralized setting simulating two entities sequentially planting backdoors into the same ResNet model on the CIFAR-10 dataset but with different target labels. Specifically, the first backdoor uses a WaNet trigger targeting class 1 while the second backdoor uses a pixel-pattern trigger targeting class 0. Following [17], the batch normalization statistics are also switched accordingly. As shown in Figure 5, the left subfigure demonstrates that the ASR of the first backdoor steadily decreases as the second backdoor is progressively injected. In contrast, the right subfigure shows that the ASR decline from normal fine-tuning (due to knowledge forgetting) remains clearly less severe, highlighting the distinct collision effect between the two backdoors.
@@
Collision Effect vs. Coexistence Effect. To this end, we can obtain a complete understanding of multi-backdoor interplay under sequential implantation: Backdoors targeting the same class tend to coexist, while those with different targets often interfere with each other. Though seemingly opposite, the two mechanisms can be understood through a unified and intuitive perspective. As noted by Indicator [17], backdoor triggers can be interpreted as OOD features in the benign latent space, typically residing in a shared OOD region within the latent space. When multiple triggers share the same target label, they construct consistent mappings from the latent OOD region to that label. This alignment allows backdoors to reinforce each other, resulting in their coexistence. In contrast, when triggers are associated with different target labels, the newly implanted trigger disrupts the existing mapping from the latent OOD region to the prior label, leading to collision or substitution.
@@
Insights from Collision Effect. Based on the collision effect, it becomes feasible to design a backdoor-based watermark that intentionally interferes with the attacker’s trigger while remaining intact in benign clients. This leads to a new proactive detection paradigm distinct from the existing coexistence-based methods, where client models that fail to retain the watermark ($i.e.$, exhibit low inspection accuracy) are identified as malicious, while those that preserve it ($i.e.$, show high inspection accuracy) are considered benign. Under this inverted paradigm, the high inspection accuracy induced by OOD prediction bias in benign clients no longer misleads detection; instead, it becomes beneficial in some cases. Building on these insights, we propose our collision-based watermarking method, as detailed in the next section.
@@
Figure 5: The collision effect between backdoors with different targets. (Left): A subsequently injected backdoor significantly degrades the ASR of the first backdoor. (Right): The ASR degradation of the first backdoor caused by standard fine-tuning is notably smaller than that caused by subsequent backdoor injection. Together, they clearly demonstrate the presence of a collision effect.
@@ 4 Methodology
Figure 6 outlines three stages of our method: watermark injection, watermark interaction, and watermark detection. We elaborate on each stage below.
@@ 4.1 Watermark Injection
Watermarking is common for integrity protection by embedding identifiable patterns into models or data [21]. Inspired by this idea and our observation of the backdoor collision effect, we leverage a backdoor-based watermarking strategy to identify malicious clients in FL. To adapt the idea of backdoor watermarking to the out-of-distribution setting, we first revisit the conceptual structure of traditional backdoor mechanisms. In conventional in-distribution scenarios, a backdoor can be viewed as composed of two parts: (1) a standard task mapping that aligns clean inputs with their ground-truth labels, and (2) a tampered mapping that redirects triggered inputs to attacker-specified labels. Following this intuition, we borrow a similar two-part structure in the OOD context: an OOD base mapping for clean OOD samples and an OOD watermark mapping for triggered ones.
@@
Planting a Regulated Base Mapping. In FL, the OOD planting set and client training data share the same global model, thus operating within the same prediction space. However, their ground-truth semantic spaces are inherently different. This misalignment introduces prediction ambiguity, where the same output value may imply conflicting meanings. For instance, label “0” might refer to the digit “0” in an EMNIST-based OOD planting set, but to “airplane” in a CIFAR-10 client task. Such ambiguity can lead to two negative effects: (1) the OOD mapping may distort client-learned semantics, reducing benign accuracy; (2) the non-i.i.d. nature of client training may corrupt the OOD mapping. To address these problems, we manually assign a fixed one-to-one label mapping $\pi_{1}$ from the OOD label space $\mathcal{Y}_{o}$ to the task label space $\mathcal{Y}_{t}$. This defines the base mapping planting:
$\pi_{1}:\mathcal{Y}_{o}\rightarrow\mathcal{Y}_{t},\quad\mathcal{D}_{p}=\{(\bm{x}_{i},\pi_{1}(y_{i}))\mid y_{i}\sim\mathcal{Y}_{o}\},$ (2)
$\mathcal{L}_{base}=\sum_{i\in\mathcal{D}_{p}}\mathrm{CE}(f(\bm{x}_{i};\bm{w}),\pi_{1}(y_{i})).$ (3)
@@
Here, $\mathcal{D}_{p}$ denotes the planting set, $\bm{w}$ represents the model parameters, and $\mathrm{CE}$ is the cross-entropy loss. Since OOD samples lack explicit semantic grounding in the task label space, the mapping’s meaning is not essential; what matters is that it remains fixed. Our choice of consistent one-to-one assignment stabilizes the ambiguity, reducing uncertainty in both OOD and task predictions. This contrasts with the design in BackdoorIndicator, which adopts random and dynamic base mappings, leading to greater instability. Instead, our method emphasizes a controlled and interpretable structure, as supported by the empirical results in Figure 7.
@@
Planting a Targeted Watermark Mapping. Building upon the established base OOD mapping, we further implant a targeted watermark mapping, analogous to the all-to-one setting [19, 8] in centralized backdoor:
$\pi_{2}:\mathcal{Y}_{o}\rightarrow y_{m},y_{m}\in\mathcal{Y}_{t},\quad\mathcal{D}_{w}=\{(\text{tri}(\bm{x}_{i},\bm{t}),\pi_{2}(y_{i}))\mid(\bm{x}_{i},y_{i})\sim\mathcal{D}_{p}\},$ (4)
@@
Following the prior work [17], we save and restore the running mean and variance of BatchNorm layers before and after watermark planting. This preserves the distinct feature distributions of the task and OOD spaces during training. In addition, to prevent the planted watermark from distorting the task-space sample-label mapping, we also apply a global model regularization term. Together, we get the overall loss: